OpenStudy (anonymous):

F(x)=-2x^2+3x-5 Solve for F(2x)

OpenStudy (anonymous):

Do you know how to plug expressions into functions? Do you know how to solve quadratic equations? Which part do you need help with?

OpenStudy (anonymous):

I did that and got -6x^2+6x-5

OpenStudy (anonymous):

Close, but not quite. \[ -2(2x)^2+3(2x)-5 = -2(2x)(2x) + 6x-5=-8x^2+6x-5\] Now, we can use the quadratic formula to find the roots. \[a=-8,\ b=6,\ c=-5\] Make sense?

OpenStudy (anonymous):

Am just not for sure if that is correct

OpenStudy (anonymous):

I meant not sure about my work not yours

OpenStudy (anonymous):

I don't follow you on the roots part.

OpenStudy (anonymous):

Hmmm, you might not need to find the roots. So maybe we are already finished. Disregard what I said.

OpenStudy (anonymous):

Oh wait never mind I follow you.

OpenStudy (anonymous):

Thank you for your help.